# Enhanced Phase Estimation via Photon-Added Two-Mode Squeezed States and Kerr Nonlinearity

### v1

@@ Acknowledgements.
$A_{1}$ $=\frac{\sqrt{2}}{2}\left(\sqrt{1-l}e^{-i\phi}x_{1}+iy_{1}\right),$
$A_{2}$ $=\frac{\sqrt{2}}{2}\left(y_{2}-i\sqrt{1-l}e^{i\phi}x_{2}\right),$
$A_{3}$ $=\frac{\sqrt{2}}{2}\left(y_{1}+i\sqrt{1-l}e^{-i\phi}x_{1}\right),$
$A_{4}$ $=\frac{\sqrt{2}}{2}\left(\sqrt{1-l}e^{i\phi}x_{2}-iy_{2}\right).$ (B7)
APPENDIX C : THE SPECIFIC DERIVATIONS AND EXPRESSIONS FOR $\left\langle I_{D}\right\rangle$ AND $\left\langle I_{D}^{2}\right\rangle$ UNDER $k=2$ CONDITION
With respect to operators that are partially obtained from the transformation relation of BS2, such as $a^{\dagger}b$, it is necessary to utilize the transformation relation of the Kerr nonlinear phase shift (Eq. (8)) and employ the IWOP technique to facilitate the calculation of the average values $\left\langle I_{D}\right\rangle$ and $\left\langle I_{D}^{2}\right\rangle$. Hence, we construct the general formula
$D_{2}\left(x\right)$
$=\left.{}_{f}\left\langle 0\right|\left\langle in\right|B_{1}^{\dagger}B_{f}^{\dagger}e^{-i\phi\left(a^{\dagger}a\right)^{2}}\left(a^{\dagger}b\right)^{x}e^{i\phi\left(a^{\dagger}a\right)^{2}}B_{f}B_{1}\left|in\right\rangle\left|0\right\rangle_{f}\right$.
$=\left\langle\psi\right|e^{-x^{2}i\phi}a^{\dagger x}e^{-2xi\phi a^{\dagger}a}b^{x}\left|\psi\right\rangle$
$=\frac{\partial^{x}}{\partial s^{x}}\left.\left\langle\psi\right|e^{-x^{2}i\phi}\exp\left[a^{\dagger}bs\right]e^{-2xi\phi a^{\dagger}a}\left|\psi\right\rangle\right|_{s=0},$ (C1)
where $x=1,2$ and for the convenience of using IWOP technology, $B_{f}B_{1}\left|in\right\rangle\left|0\right\rangle_{f}$ is converted into a quantum state that represents a coherent state, denoted as $\left|\psi\right\rangle$. Using the following normal ordering form
$e^{-2xi\phi a^{\dagger}a}=\colon\exp\left[\left(e^{-2xi\phi}-1\right)a^{\dagger}a\right]\colon$, (C2)
@@
$\int\frac{d^{2}z}{\pi}e^{\zeta\left|z\right|^{2}+\xi z+\eta z^{\ast}+fz^{2}+gz^{\ast 2}}=\frac{e^{\frac{-\zeta\xi\eta+\xi^{2}g+\eta^{2}f}{\zeta^{2}-4fg}}}{\sqrt{\zeta^{2}-4fg}},$ (C3)
we can compute $D_{2}\left(x\right)$ as
$D_{2}\left(x\right)=\frac{\begin{array}[]{c}\exp[-x^{2}i\phi-\alpha^{2}+M_{0}\\ +\frac{-M_{1}M_{2}M_{3}+M_{4}M_{2}^{2}+M_{5}M_{3}^{2}}{M_{1}^{2}-4M_{4}M_{5}}]\end{array}}{\cosh^{2}rN_{m,n}\sqrt{M_{1}^{2}-4M_{4}M_{5}}},$ (C4)
@@
$M_{0}$ $=V_{5}(-i\tanh r\left(V_{1}+V_{2}\right)\left(V_{1}-V_{2}\right)V_{5}$
$-i\left(V_{1}-V_{2}\right)\tau_{1}+\left(V_{1}+V_{2}\right)V_{4}),$
$M_{1}$ $=\tanh^{2}r\frac{4V_{1}^{2}+1+\left(1-l\right)s^{2}}{2}-1,$
$M_{2}$ $=\tau_{2}+\tanh r(\left(V_{1}+V_{2}\right)V_{4}-i\left(V_{1}-V_{2}\right)\tau_{1})$
$-i\tanh^{2}r\left(V_{1}+V_{2}\right)\left(V_{1}-V_{2}\right)V_{5},$
$M_{3}$ $=i\left(V_{1}-V_{3}\right)V_{4}+\left(V_{1}+V_{3}\right)\tau_{1}$
$+\tanh rV_{5}\frac{4V_{1}^{2}+1+\left(1-l\right)s^{2}}{2},$
$M_{4}$ $=i\tanh r\left(V_{1}+V_{3}\right)\left(V_{1}-V_{3}\right),$
$M_{5}$ $=-i\tanh^{3}r\left(V_{1}+V_{2}\right)\left(V_{1}-V_{2}\right),$ (C5)
@@
By further utilizing Eqs. (6), (B1) and (C1), the expressions for $\left\langle I_{D}\right\rangle$ and $\left\langle I_{D}^{2}\right\rangle$ can be obtained for $k=2$ as
$\left\langle I_{D}\right\rangle=i\left[D_{2}\left(1\right)-D_{2}\left(1\right)^{\ast}\right],$ (C7)
@@
$\left\langle I_{D}^{2}\right\rangle$ $=D_{1}\left(0,1,0,1\right)+D_{1}\left(1,0,1,0\right)$
$+2D_{1}\left(1,1,1,1\right)-D_{2}\left(2\right)-D_{2}\left(2\right)^{\ast}.$ (C8)
APPENDIX D : $C_{Q}$ FOR $k=2$ CASE
In this appendix, we derive $C_{Q}$ for the Kerr nonlinear phase shift with $k=2$ and its specific expression to obtain the QFI  $F_{L_{2}}$ under photon losses condition. To compute $C_{Q}$ using Eqs. (21), (22), (25) and (26) according to the method in Refs. 40 and 49 , we first obtain the normal ordering form of $\left(1-l\right)^{n_{b}}n_{a}^{q}$ by utilizing the operator identity from Eq. (16) as follows
$\left(1-l\right)^{n_{b}}n_{a}^{q}$ $=\eta^{n_{a}}n_{a}^{q}$
$=\frac{\partial^{q}}{\partial x^{q}}\left.\exp\left[n_{a}\ln\eta\right]\exp\left[n_{a}x\right]\right|_{x=0}$
$=\colon\left.\frac{\partial^{q}}{\partial x^{q}}e^{\left(\eta e^{x}-1\right)a^{\dagger}a}\right|_{x=0}\colon$, (D1)
where for simplicity, we set $\eta=1-l$. Based on this equation and using IWOP technology, the following summation expression $S_{q,p}$ can be further calculated Based on this equation and further utilizing the IWOP technique, the following summation can be computed for the operators associated with the Hermitian operators $H_{1,2}$ from Eqs. (21) and (22).
$S_{q,p}$ $=\sum_{j=0}^{\infty}\frac{\left(1-\eta\right)^{j}}{j!}j^{p}a^{\dagger j}\eta^{n_{a}}n_{a}^{q}a^{j}$
$=\sum_{j=0}^{\infty}\frac{\left(1-\eta\right)^{j}}{j!}j^{p}\colon\left.\left(a^{\dagger}a\right)^{j}\frac{\partial^{q}}{\partial x^{q}}e^{\left(\eta e^{x}-1\right)a^{\dagger}a}\right|_{x=0}\colon$
$=\colon\sum_{j=0}^{\infty}\frac{\left[\left(1-\eta\right)a^{\dagger}a\right]^{j}}{j!}\frac{\partial^{q+p}}{\partial x^{q}\partial y^{p}}\left.e^{\left(\eta e^{x}-1\right)a^{\dagger}a+yj}\right|_{x=y=0}\colon$
$=\colon\frac{\partial^{q+p}}{\partial x^{q}\partial y^{p}}\left.e^{\left[\eta e^{x}+\left(1-\eta\right)e^{y}-1\right]a^{\dagger}a}\right|_{x=y=0}\colon$
$=\frac{\partial^{q+p}}{\partial x^{q}\partial y^{p}}\left.\left[\eta e^{x}+\left(1-\eta\right)e^{y}\right]^{n_{a}}\right|_{x=y=0}.$ (D2)
@@
Similar to the calculation method for $C_{Q}$ in Eq. (20), substituting the Kraus operator with $k=2$ (Eq. (25)) into Eqs (21) and (22), and combining with Eq. (D2) and the following transformation relationships,
@@
$e^{\lambda\left(a^{\dagger}a\right)^{2}}a^{j}e^{-\lambda\left(a^{\dagger}a\right)^{2}}$ $=e^{\lambda j^{2}}a^{j}e^{-2\lambda j\left(a^{\dagger}a\right)},$ (D3)
@@
$C_{Q}$ $=4[K_{1}^{2}\left\langle\Delta^{2}n_{a}^{2}\right\rangle_{S}-K_{2}\left\langle n_{a}^{3}\right\rangle_{S}+K_{3}\left\langle n_{a}^{2}\right\rangle_{S}$
$-K_{4}\left\langle n_{a}\right\rangle_{S}-K_{5}\left\langle n_{a}^{2}\right\rangle_{S}\left\langle n_{a}\right\rangle_{S}-K_{6}\left\langle n_{a}\right\rangle_{S}^{2}],$ (D4)
@@
$K_{2}$ $=2\eta\left[3\omega_{1}^{2}\eta^{3}-3\omega_{3}^{2}\eta^{2}-\omega_{4}\eta+\omega_{5}\right],$
$K_{3}$ $=\eta\left[11\omega_{1}^{2}\eta^{3}-2\omega_{6}^{2}\eta^{2}+\omega_{7}\eta-4\omega_{1}\omega_{2}\right],$
$K_{4}$ $=\eta\omega_{1}^{2}\left(6\eta^{3}-12\eta^{2}+7\eta-1\right),$
$K_{5}$ $=2\left(1-\eta\right)\eta\omega_{1}K_{1},$
$K_{6}$ $=\left(1-\eta\right)^{2}\eta^{2}\omega_{1}^{2},$ (D5)
@@
$\omega_{3}$ $=1+2\left(3\mu_{1}-2\mu_{2}\right)+\left(2\mu_{1}-\mu_{2}\right)\left(4\mu_{1}-3\mu_{2}\right),$
@@
where parameters $\mu_{1}$ and $\mu_{2}$ are optimizable to describe the photon losses occurring before and after the phase shifter. In particular, $\mu_{1}=\mu_{2}=0$ or 1-1 represents photon losses occurring before or after the phase shifter, respectively, as depicted in Fig. 8. Specifically, for the case $\mu_{1}=\mu_{2}=-1$, $F_{L_{2}}\leqslant C_{Q}=4\left\langle\Delta^{2}n_{a}^{2}\right\rangle$, representing the ideal case of QFI. Using Eqs. (D4)-(D6), optimizing through $\partial C_{Q}/\partial\mu_{1}=\partial C_{Q}/\partial\mu_{2}=0$ to find the minimum value of $C_{Q}$ gives
$\mu_{1opt}$ $=\frac{G_{2}G_{5}-G_{3}G_{4}}{G_{1}G_{4}-2\eta G_{2}^{2}},$ (D7)
$\mu_{2opt}$ $=\frac{G_{1}G_{5}-2\eta G_{2}G_{3}}{G_{1}G_{4}-2\eta G_{2}^{2}},$ (D8)
@@
$G_{1}$ $=2\left[-\left(1-\eta\right)\eta\left(\left\langle\Delta^{2}n_{a}^{2}\right\rangle+2\left\langle n_{a}^{2}\right\rangle\left\langle n_{a}\right\rangle-\left\langle n_{a}\right\rangle^{2}\right)\right$.
$-\left(6\eta^{2}-6\eta+1\right)\left(\left\langle n_{a}^{3}\right\rangle+\left\langle n_{a}\right\rangle\right)$
$\left.+\left(11\eta^{2}-11\eta+2\right)\left\langle n_{a}^{2}\right\rangle\right],$
$G_{2}$ $=\left(1-\eta\right)^{2}\left\langle\Delta^{2}n_{a}^{2}\right\rangle+3\left(1-\eta\right)\left(2\eta-1\right)\left\langle n_{a}^{3}\right\rangle$
$+\left(11\eta^{2}-13\eta+3\right)\left\langle n_{a}^{2}\right\rangle-\left(6\eta^{2}-6\eta+1\right)\left\langle n_{a}\right\rangle$
$-\left(1-\eta\right)\left(2\eta-1\right)\left\langle n_{a}^{2}\right\rangle\left\langle n_{a}\right\rangle+\eta\left(1-\eta\right)\left\langle n_{a}\right\rangle^{2},$ (D9)
@@
$G_{3}$ $=\eta^{2}\left\langle\Delta^{2}n_{a}^{2}\right\rangle-3\eta\left(2\eta-1\right)\left\langle n_{a}^{3}\right\rangle$
$+\left(11\eta^{2}-9\eta+1\right)\left\langle n_{a}^{2}\right\rangle-\left(6\eta^{2}-6\eta+1\right)\left\langle n_{a}\right\rangle$
$+\eta\left(2\eta-1\right)\left\langle n_{a}^{2}\right\rangle\left\langle n_{a}\right\rangle+\eta\left(1-\eta\right)\left\langle n_{a}\right\rangle^{2},$
$G_{4}$ $=-\left(1-\eta\right)^{3}\left\langle\Delta^{2}n_{a}^{2}\right\rangle-6\eta\left(1-\eta\right)^{2}\left\langle n_{a}^{3}\right\rangle$
$-\eta\left(1-\eta\right)\left(11\eta-4\right)\left\langle n_{a}^{2}\right\rangle-\eta\left(6\eta^{2}-6\eta+1\right)\left\langle n_{a}\right\rangle$
$+2\eta\left(1-\eta\right)^{2}\left\langle n_{a}^{2}\right\rangle\left\langle n_{a}\right\rangle+\eta^{2}\left(1-\eta\right)\left\langle n_{a}\right\rangle^{2},$
$G_{5}$ $=\eta\left[-\eta\left(1-\eta\right)\left(\left\langle\Delta^{2}n_{a}^{2}\right\rangle-\left\langle n_{a}\right\rangle^{2}\right)\right$.
$-\left(6\eta^{2}-6\eta+1\right)\left(\left\langle n_{a}^{3}\right\rangle+\left\langle n_{a}\right\rangle\right)$
$+\left(11\eta^{2}-11\eta+2\right)\left\langle n_{a}^{2}\right\rangle$
$\left.+\left(2\eta^{2}-2\eta+1\right)\left\langle n_{a}^{2}\right\rangle\left\langle n_{a}\right\rangle\right],$ (D10)

### v2

@@ Appendix A Derivation of Transformation Relation for Kerr Nonlinear Phase Shifter Operator
$A_{1}$ $=\frac{\sqrt{2}}{2}\bigl(\sqrt{1-l}e^{-i\phi}x_{1}+iy_{1}\bigr),$
$A_{2}$ $=\frac{\sqrt{2}}{2}\bigl(y_{2}-i\sqrt{1-l}e^{i\phi}x_{2}\bigr),$
$A_{3}$ $=\frac{\sqrt{2}}{2}\bigl(y_{1}+i\sqrt{1-l}e^{-i\phi}x_{1}\bigr),$
$A_{4}$ $=\frac{\sqrt{2}}{2}\bigl(\sqrt{1-l}e^{i\phi}x_{2}-iy_{2}\bigr).$ (B7)
Appendix C Specific Derivations and Expressions for $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ under $k=2$ Condition
With respect to operators that are partially obtained from the transformation relation of BS2, such as $a^{\dagger}b$, it is necessary to utilize the transformation relation of the Kerr nonlinear phase shift (Eq. (8)) and employ the IWOP technique to facilitate the calculation of the average values $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$. Hence, we construct the general formula
$D_{2}(x)$
$={}_{f}\langle 0\rvert\langle\text{in}\rvert B_{1}^{\dagger}B_{f}^{\dagger}e^{-i\phi(a^{\dagger}a)^{2}}(a^{\dagger}b)^{x}e^{i\phi(a^{\dagger}a)^{2}}B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}$
$=\langle\psi\rvert e^{-x^{2}i\phi}a^{\dagger x}e^{-2xi\phi a^{\dagger}a}b^{x}\lvert\psi\rangle$
$=\frac{\partial^{x}}{\partial s^{x}}\langle\psi\rvert e^{-x^{2}i\phi}\exp[a^{\dagger}bs]e^{-2xi\phi a^{\dagger}a}\lvert\psi\rangle\Big|_{s=0},$ (C1)
where $x=1,2$ and for the convenience of using IWOP technology, $B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}$ is converted into a quantum state that represents a coherent state, denoted as $\lvert\psi\rangle$. Using the following normal ordering form
$e^{-2xi\phi a^{\dagger}a}=\colon\!\exp\bigl[(e^{-2xi\phi}-1)a^{\dagger}a\bigr]\!\colon$, (C2)
@@
$\int\frac{d^{2}z}{\pi}e^{\zeta\lvert z\rvert^{2}+\xi z+\eta z^{\ast}+fz^{2}+gz^{\ast 2}}=\frac{e^{\frac{-\zeta\xi\eta+\xi^{2}g+\eta^{2}f}{\zeta^{2}-4fg}}}{\sqrt{\zeta^{2}-4fg}},$ (C3)
we can compute $D_{2}(x)$ as
$D_{2}(x)=\frac{\exp\bigl[-x^{2}i\phi-\alpha^{2}+M_{0}+\frac{-M_{1}M_{2}M_{3}+M_{4}M_{2}^{2}+M_{5}M_{3}^{2}}{M_{1}^{2}-4M_{4}M_{5}}\bigr]}{\cosh^{2}r\,N_{m,n}\sqrt{M_{1}^{2}-4M_{4}M_{5}}},$ (C4)
@@
$M_{0}$ $=V_{5}\bigl(-i\tanh r(V_{1}+V_{2})(V_{1}-V_{2})V_{5}$
$\quad-i(V_{1}-V_{2})\tau_{1}+(V_{1}+V_{2})V_{4}\bigr),$
$M_{1}$ $=\tanh^{2}r\frac{4V_{1}^{2}+1+(1-l)s^{2}}{2}-1,$
$M_{2}$ $=\tau_{2}+\tanh r\bigl((V_{1}+V_{2})V_{4}-i(V_{1}-V_{2})\tau_{1}\bigr)$
$\quad-i\tanh^{2}r(V_{1}+V_{2})(V_{1}-V_{2})V_{5},$
$M_{3}$ $=i(V_{1}-V_{3})V_{4}+(V_{1}+V_{3})\tau_{1}$
$\quad+\tanh rV_{5}\frac{4V_{1}^{2}+1+(1-l)s^{2}}{2},$
$M_{4}$ $=i\tanh r(V_{1}+V_{3})(V_{1}-V_{3}),$
$M_{5}$ $=-i\tanh^{3}r(V_{1}+V_{2})(V_{1}-V_{2}),$ (C5)
@@
By further utilizing Eqs. (6), (B1) and (C1), the expressions for $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ can be obtained for $k=2$ as
$\langle I_{D}\rangle=i\bigl[D_{2}(1)-D_{2}(1)^{\ast}\bigr],$ (C7)
@@
$\langle I_{D}^{2}\rangle$ $=D_{1}(0,1,0,1)+D_{1}(1,0,1,0)$
$\quad+2D_{1}(1,1,1,1)-D_{2}(2)-D_{2}(2)^{\ast}.$ (C8)
Appendix D $C_{Q}$ for $k=2$ Case
In this appendix, we derive $C_{Q}$ for the Kerr nonlinear phase shift with $k=2$ and its specific expression to obtain the QFI $F_{L_{2}}$ under photon losses condition. To compute $C_{Q}$ using Eqs. (21), (22), (25) and (26) according to the method in Refs. 40 and 49 , we first obtain the normal ordering form of $(1-l)^{n_{a}}n_{a}^{q}$ by utilizing the operator identity from Eq. (16) as follows
$(1-l)^{n_{a}}n_{a}^{q}$ $=\eta^{n_{a}}n_{a}^{q}$
$=\frac{\partial^{q}}{\partial x^{q}}\exp\bigl[n_{a}\ln\eta\bigr]\exp\bigl[n_{a}x\bigr]\Big|_{x=0}$
$=\colon\!\frac{\partial^{q}}{\partial x^{q}}e^{(\eta e^{x}-1)a^{\dagger}a}\Big|_{x=0}\!\colon$, (D1)
where for simplicity, we set $\eta=1-l$. Based on this equation and using IWOP technology, the following summation expression $S_{q,p}$ can be further calculated for the operators associated with the Hermitian operators $H_{1,2}$ from Eqs. (21) and (22):
$S_{q,p}$ $=\sum_{j=0}^{\infty}\frac{(1-\eta)^{j}}{j!}j^{p}a^{\dagger j}\eta^{n_{a}}n_{a}^{q}a^{j}$
$=\sum_{j=0}^{\infty}\frac{(1-\eta)^{j}}{j!}j^{p}\colon\!(a^{\dagger}a)^{j}\frac{\partial^{q}}{\partial x^{q}}e^{(\eta e^{x}-1)a^{\dagger}a}\Big|_{x=0}\!\colon$
$=\colon\!\sum_{j=0}^{\infty}\frac{[(1-\eta)a^{\dagger}a]^{j}}{j!}\frac{\partial^{q+p}}{\partial x^{q}\partial y^{p}}e^{(\eta e^{x}-1)a^{\dagger}a+yj}\Big|_{x=y=0}\!\colon$
$=\colon\!\frac{\partial^{q+p}}{\partial x^{q}\partial y^{p}}e^{[\eta e^{x}+(1-\eta)e^{y}-1]a^{\dagger}a}\Big|_{x=y=0}\!\colon$
$=\frac{\partial^{q+p}}{\partial x^{q}\partial y^{p}}\bigl[\eta e^{x}+(1-\eta)e^{y}\bigr]^{n_{a}}\Big|_{x=y=0}.$ (D2)
@@ Appendix D $C_{Q}$ for $k=2$ Case
Similar to the calculation method for $C_{Q}$ in Eq. (20), substituting the Kraus operator with $k=2$ (Eq. (25)) into Eqs. (21) and (22), and combining with Eq. (D2) and the following transformation relationships,
@@
$e^{\lambda(a^{\dagger}a)^{2}}a^{j}e^{-\lambda(a^{\dagger}a)^{2}}$ $=e^{\lambda j^{2}}a^{j}e^{-2\lambda ja^{\dagger}a},$ (D3)
@@
$C_{Q}$ $=4\bigl[K_{1}^{2}\langle\Delta^{2}n_{a}^{2}\rangle_{S}-K_{2}\langle n_{a}^{3}\rangle_{S}+K_{3}\langle n_{a}^{2}\rangle_{S}$
$\quad-K_{4}\langle n_{a}\rangle_{S}-K_{5}\langle n_{a}^{2}\rangle_{S}\langle n_{a}\rangle_{S}-K_{6}\langle n_{a}\rangle_{S}^{2}\bigr],$ (D4)
@@
$K_{2}$ $=2\eta\bigl[3\omega_{1}^{2}\eta^{3}-3\omega_{3}^{2}\eta^{2}-\omega_{4}\eta+\omega_{5}\bigr],$
$K_{3}$ $=\eta\bigl[11\omega_{1}^{2}\eta^{3}-2\omega_{6}^{2}\eta^{2}+\omega_{7}\eta-4\omega_{1}\omega_{2}\bigr],$
$K_{4}$ $=\eta\omega_{1}^{2}(6\eta^{3}-12\eta^{2}+7\eta-1),$
$K_{5}$ $=2(1-\eta)\eta\omega_{1}K_{1},$
$K_{6}$ $=(1-\eta)^{2}\eta^{2}\omega_{1}^{2},$ (D5)
@@
$\omega_{3}$ $=1+2(3\mu_{1}-2\mu_{2})+(2\mu_{1}-\mu_{2})(4\mu_{1}-3\mu_{2}),$
@@
where parameters $\mu_{1}$ and $\mu_{2}$ are optimizable to describe the photon losses occurring before and after the phase shifter. In particular, $\mu_{1}=\mu_{2}=0$ or 1-1 represents photon losses occurring before or after the phase shifter, respectively, as depicted in Fig. 7. Specifically, for the case $\mu_{1}=\mu_{2}=-1$, $F_{L_{2}}\leqslant C_{Q}=4\langle\Delta^{2}n_{a}^{2}\rangle$, representing the ideal case of QFI. Using Eqs. (D4)–(D6), optimizing through $\partial C_{Q}/\partial\mu_{1}=\partial C_{Q}/\partial\mu_{2}=0$ to find the minimum value of $C_{Q}$ gives
$\mu_{1\text{opt}}$ $=\frac{G_{2}G_{5}-G_{3}G_{4}}{G_{1}G_{4}-2\eta G_{2}^{2}},$ (D7)
$\mu_{2\text{opt}}$ $=\frac{G_{1}G_{5}-2\eta G_{2}G_{3}}{G_{1}G_{4}-2\eta G_{2}^{2}},$ (D8)
@@
$G_{1}$ $=2\bigl[-(1-\eta)\eta\bigl(\langle\Delta^{2}n_{a}^{2}\rangle+2\langle n_{a}^{2}\rangle\langle n_{a}\rangle-\langle n_{a}\rangle^{2}\bigr)$
$\quad-(6\eta^{2}-6\eta+1)\bigl(\langle n_{a}^{3}\rangle+\langle n_{a}\rangle\bigr)$
$\quad+(11\eta^{2}-11\eta+2)\langle n_{a}^{2}\rangle\bigr],$
$G_{2}$ $=(1-\eta)^{2}\langle\Delta^{2}n_{a}^{2}\rangle+3(1-\eta)(2\eta-1)\langle n_{a}^{3}\rangle$
$\quad+(11\eta^{2}-13\eta+3)\langle n_{a}^{2}\rangle-(6\eta^{2}-6\eta+1)\langle n_{a}\rangle$
$\quad-(1-\eta)(2\eta-1)\langle n_{a}^{2}\rangle\langle n_{a}\rangle+\eta(1-\eta)\langle n_{a}\rangle^{2},$ (D9)
@@
$G_{3}$ $=\eta^{2}\langle\Delta^{2}n_{a}^{2}\rangle-3\eta(2\eta-1)\langle n_{a}^{3}\rangle$
$\quad+(11\eta^{2}-9\eta+1)\langle n_{a}^{2}\rangle-(6\eta^{2}-6\eta+1)\langle n_{a}\rangle$
$\quad+\eta(2\eta-1)\langle n_{a}^{2}\rangle\langle n_{a}\rangle+\eta(1-\eta)\langle n_{a}\rangle^{2},$
$G_{4}$ $=-(1-\eta)^{3}\langle\Delta^{2}n_{a}^{2}\rangle-6\eta(1-\eta)^{2}\langle n_{a}^{3}\rangle$
$\quad-\eta(1-\eta)(11\eta-4)\langle n_{a}^{2}\rangle-\eta(6\eta^{2}-6\eta+1)\langle n_{a}\rangle$
$\quad+2\eta(1-\eta)^{2}\langle n_{a}^{2}\rangle\langle n_{a}\rangle+\eta^{2}(1-\eta)\langle n_{a}\rangle^{2},$
$G_{5}$ $=\eta\bigl[-\eta(1-\eta)\bigl(\langle\Delta^{2}n_{a}^{2}\rangle-\langle n_{a}\rangle^{2}\bigr)$
$\quad-(6\eta^{2}-6\eta+1)\bigl(\langle n_{a}^{3}\rangle+\langle n_{a}\rangle\bigr)$
$\quad+(11\eta^{2}-11\eta+2)\langle n_{a}^{2}\rangle$
$\quad+(2\eta^{2}-2\eta+1)\langle n_{a}^{2}\rangle\langle n_{a}\rangle\bigr],$ (D10)